OpenStudy (devonhoward15):

which equation is the direct linear variation equation for the relationship? the variable y varies directly with x and y=12 when x=3... will fan and medal!!! A.y=x+9 B.y=x-9 C.y=4x D.y=1/4x

OpenStudy (anonymous):

What do you think it is?

OpenStudy (devonhoward15):

a

OpenStudy (devonhoward15):

cause if you replace x with 3 it equals 12 which is equal to y

OpenStudy (anonymous):

Correct

OpenStudy (devonhoward15):

thanks

OpenStudy (devonhoward15):

are you able to help me with 2 more questions?

OpenStudy (anonymous):

Sure

OpenStudy (devonhoward15):

ok thanks

OpenStudy (devonhoward15):

what is the inverse variation equation for the relationship? y varies inversely with x and y =4 when x=2

OpenStudy (devonhoward15):

A.y=2x B.y=x+2 C.y=1/2x D.y=8/x

OpenStudy (anonymous):

A because put numbers in 4y=2(2x) 2 times to = 4x so 4y=4x ":)

OpenStudy (devonhoward15):

thank you... got 1 more if you dont mind

OpenStudy (anonymous):

sure

OpenStudy (devonhoward15):

the function f(x) varies directly with x and f(x)=50 when x=-10 what id f(x) when x=15

OpenStudy (devonhoward15):

A.-30 B.-75 C.-150 D.-5

OpenStudy (anonymous):

i would say B?

OpenStudy (devonhoward15):

ok.... thanks very much

OpenStudy (devonhoward15):

well im going to bed goodnight